# Hermitian and non-Hermitian topology in active matter

### v1

@@ 5.4 Real-space topology
Despite the abundance of studies of topological defects in various fields of physics, those found in active matter have been mainly restricted to defects in two dimensions [35, 70, 71, 72]. Thus, exploring topological defects in, e.g., three-dimensional active matter can enrich the interplay between the real-space topology and active matter dynamics. In addition, the Kibble-Zurek mechanism may bring new insight into the analysis of the dynamics of topological defects. Since topological defects can induce effective boundaries and emergent gauge fields, the nontrivial interplay between the band topology and topological defects may enrich the topological phenomena in active matter.
@@ 5.5 Biological functions
So far, there has been no example of a topological edge mode at play in real biological tissues. In cultured cells, bioengineering of edge dynamics at the interface between tissues has been implemented in Refs. [279, 315]. While the topological origin of these edge dynamics is still under investigation, their existence implies the possibility that topological edge modes can work in real biological systems. In a broader perspective, some characteristics of topological edge modes, such as localization [32], chiral motion [316], and robustness [105] might be found throughout biology. For example, robustness should be important for biological cells to adjust the changing environments. It will be interesting if the studies of topological active matter uncover the hidden topological structures in biology and elucidate the mechanism of, e.g., the robustness in biological functions.
We are grateful to Zongping Gong, Daiki Nishiguchi, Takahiro Sagawa, Taro Sawada, and Tsuneya Yoshida for fruitful discussions. K.S. acknowledges support from JSPS KAKENHI Grant No. JP21J20199 and World-leading Innovative Graduate Study Program for Materials Research, Information, and Technology (MERIT-WINGS) of the University of Tokyo. K.Y. acknowledges support from JSPS KAKENHI through Grant No. JP21J01409. K.K. acknowledges support from JSPS KAKENHI Grants No. JP19H05795, No. JP19H05275, No. JP21H01007, and No. JP23H00095. Y.A. acknowledges support from the Japan Society for the Promotion of Science (JSPS) through Grant No. JP19K23424 and from JST FOREST Program (Grant Number JPMJFR [^2] U, Japan) and JST CREST (Grant Number JPMJCR23I2, Japan).

### v2

@@ 6.4 Real-space topology
Despite the abundance of studies of topological defects in various fields of physics, those found in active matter have been mainly restricted to defects in two dimensions [33, 73, 74, 75]. Thus, exploring topological defects in, e.g., three-dimensional active matter [322] can enrich the interplay between the real-space topology and active matter dynamics. In addition, the Kibble-Zurek mechanism may bring new insight into the analysis of the dynamics of topological defects. Since topological defects can induce effective boundaries and emergent gauge fields, the nontrivial interplay between the band topology and topological defects may enrich the topological phenomena in active matter.
@@ 6.5 Biological functions
So far, there has been no example of a topological edge mode at play in real biological systems (i.e., in multicellular tissues, inside cells, or even in purified biomolecular systems in vitro). In cultured cells, bioengineering of edge dynamics at the interface between tissues has been implemented in Refs. [287, 331]. While the topological origin of these edge dynamics is still under investigation, their existence implies the possibility that topological edge modes can work in real biological systems. In a broader perspective, some characteristics of topological edge modes, such as localization [30], chiral motion [332], and robustness [108] might be found throughout biology. For example, robustness should be important for biological cells to adjust to changing environments. It will be interesting if the studies of topological active matter uncover the hidden topological structures in biology and elucidate the mechanism of, e.g., the robustness in biological functions.
We are grateful to Zongping Gong, Daiki Nishiguchi, Takahiro Sagawa, Taro Sawada, and Tsuneya Yoshida for fruitful discussions. K.S. acknowledges support from JSPS KAKENHI Grant No. JP21J20199, JP24K22848, and World-leading Innovative Graduate Study Program for Materials Research, Information, and Technology (MERIT-WINGS) of the University of Tokyo. K.Y. acknowledges support from JSPS KAKENHI through Grant No. JP21J01409. $K.$ acknowledges support from JSPS KAKENHI Grants No. JP19H05795, No. JP19H05275, No. JP21H01007, and No. JP23H00095. Y.A. acknowledges support from the Japan Society for the Promotion of Science (JSPS) through Grant No. JP19K23424 and from JST FOREST Program (Grant Number JPMJFR [^2] U, Japan) and JST CREST (Grant Number JPMJCR23I2, Japan).